Sue's average for 9 games of bowling is 108. What is the lowest score she can receive for the tenth game to have an average of 110?
Answer by checkley77(12569)

9*108+x=10*110
972+x=1100
x=1100-972
x=128 must be the tenth game to average 110.
9*108+128=1100
972+128=1100
1100=1100